OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (0,-6) and (3,15)

OpenStudy (anonymous):

\[Slope = m=\frac{ y2-y1 }{ x2-x1 } Slope intercept form= y-y1=m(x-x1) \]

OpenStudy (anonymous):

\[slope= m= \frac{ 15-(-6) }{ 3-0 }=7 slope intercept form= y-15=7(x-3) \]

OpenStudy (anonymous):

So, now 7 x (-3)= -21 and 7 x "x" = 7x then y-15=7x-21 add the 15 to both sides +15 +15 Thus, the equation equals y=7x-6